OpenStudy (anonymous):

HEEELP! Factor. x^2 -(y - z)^2

OpenStudy (anonymous):

Complete the square. Here's a property. \[a^2-(b-1)^2=[a-(b-1)][a+(b-1)]\]

OpenStudy (anonymous):

Difference between two squares sorry.

OpenStudy (anonymous):

It's not completing the square. I got it mixed up. It's difference between two squares!

OpenStudy (anonymous):

It's the same as \[x^2-1=(x-1)(x+1)\] Do you get me @peekaboopork

OpenStudy (anonymous):

Yes I do, thank you. It makes so much sense now

OpenStudy (anonymous):

YOu're welcome.